a hockey team played 72 games during the regular season. they won 6 more than twice as many as they lost. how many did they win ?
Answer by JBarnum(2146)

this is a funny looking age problem...
72 total games played
Wins, Losses
W+L=72
W=6+2L
6+2l+L=72
6+3L=72
3L=66
L=22
72-22=W
50=W